An environmental consultant is considering the installation of a water storage tank for a client. The tank is estimated to have an initial cost of $360,000, and annual maintenance costs are estimated to be $7,600 per year. As an alternative, a holding pond can be provided a short distance away at an initial cost of $182,000 for the pond plus $90,000 for pumps and piping. Annual operating and maintenance costs for the pumps and holding pond are estimated to be $10,000. The planning horizon is 20 years, and at that time, neither alternative has any salvage value. MARR is 21 %/year.
Answer #1

NPW of tank = -360000 - 7600 * (P/A, 21%,20)

= -360000 - 7600 * 4.656691

= -395390.85

NPW of pond = -(182000+90000) - 10000 * (P/A, 21%,20)

= -272000 - 10000 * 4.656691

= -318566.91

As the present cost of pond is less, it should be selected